OpenStudy (anonymous):

Mike and Jamal are 9 miles apart, and are planning to meet up. Mike is walking at an average speed of 3 miles per hour to meet Jamal. Jamal is driving at an average speed of 25 miles per hour to meet Mike.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Which equation can be used to find t, the time it takes for Mike and Jamal to meet? 25t – 3t = 0 25t – 3t = 9 25t + 3t = 1 25t + 3t = 9

OpenStudy (misty1212):

don't you hate word problems?

OpenStudy (misty1212):

and look at those answers!!

OpenStudy (anonymous):

Haha.. yeah. Sometimes they get a little difficult. But, I want to know how to do this properly so I don't mess up on my exam!

OpenStudy (misty1212):

i would have said solve \[28t=9\] or just compute \[\frac{9}{28}\] but i guess they want you do pretend you are an idiot and not know that \(25t+3t=28t\)

OpenStudy (anonymous):

Thank you! :) I got it now!